OpenStudy (anonymous):

the total of 5.68 times a number and 132.7 is the number minus 29.228

OpenStudy (anonymous):

I do not under stand how to do algebra....... Would appreciate any help - thanks

OpenStudy (cwrw238):

5.68x + 132.7 = x - 29.228 solve this for x and you've got your number

OpenStudy (anonymous):

What do I do from that point I do not understand at all

OpenStudy (cwrw238):

to solve this you first bring the terms in x to left hand side to do this you subtract x from both sides ( so x disappears from the righthand side) so 5.68x - x + 132.7 = -29.228 now subtract 132.7 from both sides so it disappears from the left side 5.68x - x = -29.228 - 132.7 now simplify both sides 4.68x = -161.928 divide both sides by 4.68 to get value of x x = -161.928 / 4.68 x = -34.6 thats the solution